Solve for y
4/9y -5 = 3
3/7y -4= 2


Answer by mukhopadhyay(490)

4/9y -5 = 3
=> (4/9)y = 8
=> y = 8(9/4) = 18
.........
3/7y -4= 2
=> (3/7)y = 6
=> y = 6(7/3) = 14